OpenStudy (anonymous):

| log(0.5)tan[PI/3] | / log(0.5)tan[PI/3] Log bases are in ( )

OpenStudy (anonymous):

1?

OpenStudy (anonymous):

\[\LARGE \frac{\log_{0.5}\tan \cfrac{\pi}{3} }{\log_{0.5}\tan\cfrac{\pi}{3}}\] ?? LOL

OpenStudy (anonymous):

Upper one is an absolute value

OpenStudy (ash2326):

\[\frac{| log_{(0.5)}tan[\pi/3] |} { log_{(0.5)}tan[\pi/3]} \] \[tan \pi/3= \sqrt 3\] One point to note that 0.5 is less than 1 so 0.5 has to be raised to a negative power to get \(\sqrt 3\) so we have \[\frac{|negative|}{negative}\] \[\huge \frac{positive}{negative}=-1\]

OpenStudy (anonymous):

Thank you! I have two more of them, I'll put them to the left.